DasiaTexannaJackson:

Tina's salary was $20,350 last year. This year she is getting paid $2,000 more. What was the percent increase to the nearest percent?

DasiaTexannaJackson:

Tina's salary was $20,350 last year. This year she is getting paid $2,000 more. What was the percent increase to the nearest percent? Select the best answer from the choices provided. 10% 15% 20% 100% These are the answer choices guys above.

Angle:

\[\large percent~increase = \frac{new~amount - old~amount}{old~amount} * 100\] so the old amount = 20350 and the new amount = 20350 + 2000 = 22350 so we should try: \[\frac{ 22350-20350 }{ 20350 } * 100 = ?\]

DasiaTexannaJackson:

thanks for the help Angle